# Limited Memory LRSGA Optimizer to competitive optimization

### v1

@@ 5.1 Spectral and Stability Metrics
The notions of stable, marginally stable, and unstable dynamics follow the standard classification adopted in the spectral analysis of differentiable and adversarial games Mescheder et al. (2018); Balduzzi et al. (2018); Fiez and Ratliff (2021). In discrete-time settings, linear stability is determined by the spectral radius of the estimated Jacobian: $\rho<1$ indicates convergence (stable regime), $\rho=1$ corresponds to marginal stability, and $\rho>1$ implies divergence. To account for numerical noise and finite-length trajectories, a tolerance band of $\varepsilon=0.05$ is applied. Cases with $\rho\in[1-\varepsilon,\,1+\varepsilon]$ are further analyzed by inspecting the frequency content of the loss trajectories. This inspection is based on the Fourier transform of the temporal loss signals, where the tolerance $\varepsilon$ effectively separates slowly varying components from high-frequency oscillatory modes. The spectral energy of the oscillatory band is estimated using Welch’s method Welch (1967), which averages modified periodograms over overlapping windows to reduce the variance of the power spectral density (PSD) estimate. If the high-frequency power is small, the dynamics are classified as marginally stable; otherwise, the regime is deemed unstable. This refinement is consistent with the interpretation of lightly damped rotational dynamics commonly observed in adversarial training systems Fiez and Ratliff (2021).
@@ 5.2 Experimental Setup
All experiments were conducted on the same MNIST GAN architecture used for the FID evaluation, ensuring full comparability of results. The training dynamics of three optimizers—Adam, LM-LRSGA, and LM-LRSGA-EMA—were analyzed under two fair learning-rate regimes:
@@
    LM-LRSGA-optimal setup: $\eta=0.15$, $\tau=0.002$, corresponding to the FID best-performing configuration for both LM-LRSGA and LM-LRSGA-EMA;
@@
A tolerance parameter of $\varepsilon=0.05$ was adopted in all experiments for the stability classification, as discussed in Section 5.1. All analyses were performed using Python 3.10 with the NumPy, SciPy, and PyTorch libraries.
@@ 5.3 Results and Discussion
Figures 4 and 5 provide a comprehensive visualization of the spectral and stability diagnostics obtained for the analyzed optimizers. The first row reports the evolution of the generator and discriminator losses shown on a logarithmic scale (left), the corresponding training stability metric (center), and the estimated spectrum of the local Jacobian in the complex plane (right). The spectral plot displays the eigenvalues computed from the block-Lanczos approximation of the empirical Jacobian, with the dashed unit circle indicating the boundary between contractive and divergent dynamics. The second row shows the evolution of the generator parameters (left) and discriminator parameters (center), represented by the Euclidean distance from their initialization, and the mode-collapse indicator (right). A nearly monotonic decrease of these distances together with a slow, smooth decay of the variance indicates stable and non-collapsing behavior. Overall, the combined diagnostics reveal that both LM-LRSGA variants maintain bounded losses, regular parameter trajectories, and eigenvalues concentrated within or near the unit circle, in contrast to the divergent and oscillatory dynamics observed with Adam.
In particular, the results reported in Table 1 show that when the learning rate was set to the LM-LRSGA-optimal configuration ($\eta=0.15$, $\tau=0.002$), Adam exhibited a spectral radius of $\rho=2.81$, indicating a strongly divergent regime characterized by high-frequency oscillations and large-amplitude loss fluctuations. The Fourier spectra estimated through Welch’s method Welch (1967) confirmed a concentration of power in the high-frequency range, consistent with unstable dynamics. In contrast, both LM-LRSGA and LM-LRSGA-EMA maintained nearly marginal stability with $\rho\approx 1.12$ and $\rho\approx 1.14$, respectively. Both optimizers exhibited almost constant global stability indices ($\approx 0.99$) and parameter oscillations of order $\mathcal{O}(10^{-4})$, with negligible generator variance drift ($\dot{\sigma}^{2}_{G}\approx 0$), confirming the absence of mode collapse and aligning with the superior FID results previously observed.
@@
Table 1: Spectral and stability metrics for Adam, LM-LRSGA, and LM-LRSGA-EMA under two learning-rate regimes. All quantities are computed from block-Lanczos spectral estimates and the auxiliary stability indices described in Section 5.1.
Setting Optimizer $\rho$ Pred. Stab. Stab. (G/D) Mode-Collapse
$\eta=0.15$, $\tau=0.002$ Adam 2.81 Unstable 0.30 / 0.72 High
LM-LRSGA 1.11 Marginal 0.99 / 0.99 Low
LM-LRSGA-EMA 1.14 Marginal 0.99 / 0.99 Low
$\eta=0.001$, $\tau=0.002$ Adam 1.48 Unstable 0.99 / 0.98 Low
LM-LRSGA 1.06 Marginal Stable 1.00 / 1.00 Low
LM-LRSGA-EMA 1.06 Marginal Stable 1.00 / 1.00 Low
@@
Across both learning-rate regimes, the limited-memory variants substantially reduce the spectral radius—by approximately 50% relative to Adam—and maintain stable trajectories even for large step sizes. The reduction of $\rho$ and the absence of high-frequency energy in the loss spectra confirm that the LM-LRSGA correction effectively damps the antisymmetric component of the gradient field, yielding dynamics that are nearly contractive in the sense of the spectral criterion. These findings provide a consistent spectral explanation for the smoother convergence and improved sample quality obtained by LM-LRSGA in the FID experiments.
@@
(c) LM-LRSGA-EMA
Figure 4: Spectral and stability diagnostics for Adam, LM-LRSGA, and LM-LRSGA-EMA with learning rate $\eta=0.15$. Each panel shows (top) loss evolution in logarithmic scale, the training stability metric, and the estimated Jacobian spectrum in the complex plane; and (bottom) the generator and discriminator parameter evolution together with the mode-collapse indicator.
@@
(c) LM-LRSGA-EMA
Figure 5: Spectral and stability diagnostics for Adam, LM-LRSGA, and LM-LRSGA-EMA with learning rate $\eta=0.001$. Each panel shows (top) loss evolution in logarithmic scale, the training stability metric, and the estimated Jacobian spectrum in the complex plane; and (bottom) the generator and discriminator parameter evolution together with the mode-collapse indicator.
@@ 6 Conclusions
This work introduced Limited-Memory Low-Rank Symplectic Gradient Adjustment (LM-LRSGA), a scalable optimization framework for differentiable games and competitive learning. Building on symplectic gradient corrections and quasi-Newton approximations, LM-LRSGA retains the stabilizing effect of second-order information while reducing memory and compute via two-loop recursions algorithms. By storing only a short history of curvature pairs, it attains near–second-order stability with first-order–like efficiency.
Our analyses show that LM-LRSGA mitigates the rotational dynamics that typically hinder convergence in adversarial settings. On GAN training with MNIST and Fashion-MNIST, it yields smoother, synchronized loss trajectories, faster convergence, and up to $30\%$ FID improvement over Adam. These results indicate that limited-memory curvature corrections offer a practical balance of stability, efficiency, and scalability for competitive optimization.
Future work will focus on consolidating the theoretical foundations. We aim to develop a more comprehensive analysis for nonconvex and locally monotone regimes, where classical tools are less informative. Such a framework would clarify the role of curvature information and symplectic corrections beyond convex–concave settings and inform adaptive limited-memory strategies for broader classes of differentiable games.
LM-LRSGA also supports sustainable, energy-aware machine learning. By avoiding explicit formation and storage of high-dimensional second-order tensors, it reduces memory transfers and expensive operations—two key drivers of energy use in large-scale training—thereby lowering runtime in line with “green AI” principles. Thus, LM-LRSGA provides a step toward methods that are not only theoretically and numerically robust, but also environmentally sustainable.

### v2

@@ 7.1 Spectral and Stability Metrics
The notions of stable, marginally stable, and unstable dynamics follow the standard classification adopted in the spectral analysis of differentiable and adversarial games Mescheder et al. (2018); Balduzzi et al. (2018); Fiez and Ratliff (2021). In discrete-time settings, linear stability is determined by the spectral radius of the estimated Jacobian: $\rho<1$ indicates convergence (stable regime), $\rho=1$ corresponds to marginal stability, and $\rho>1$ implies divergence. To account for numerical noise and finite-length trajectories, a tolerance band of $\epsilon=0.15$ is applied. Cases with $\rho\in[1-\epsilon,\,1+\epsilon]$ are further analyzed by inspecting the frequency content of the loss trajectories. This inspection is based on the Fourier transform of the temporal loss signals, where the tolerance $\epsilon$ effectively separates slowly varying components from high-frequency oscillatory modes. The spectral energy of the oscillatory band is estimated using Welch’s method Welch (1967), which averages modified periodograms over overlapping windows to reduce the variance of the power spectral density (PSD) estimate. In practice, we compute a high-frequency power index for both the generator and the discriminator; when these indices are of order $10^{0}$, the dynamics are classified as marginally stable, whereas regimes with significantly larger high-frequency power (i.e., well above $10^{0}$) are deemed unstable. This refinement is consistent with the interpretation of lightly damped rotational dynamics commonly observed in adversarial training systems Fiez and Ratliff (2021).
@@ 7.2 Experimental Setup
All experiments were conducted on the same MNIST GAN architecture used for the FID evaluation, ensuring full comparability of results. The training dynamics of the three optimizers (Adam, LM-LRSGA, and LM-LRSGAEma) were analyzed under two fair learning-rate regimes:
@@
    LM-LRSGA-optimal setup: $\eta=0.2$, $\tau=0.002$, corresponding to the FID best-performing configuration for both LM-LRSGA and LM-LRSGAEma;
@@
A tolerance parameter of $\epsilon=0.15$ was adopted in all experiments for the stability classification, as discussed in Section 7.1. All analyses were performed using Python 3.10 with the NumPy, SciPy, and PyTorch libraries.
@@ 7.3 Results and Discussion
Figures 4 and 5 provide a comprehensive visualization of the spectral and stability diagnostics obtained for the analyzed optimizers. The first row reports the evolution of the generator and discriminator losses shown on a logarithmic scale (left), the corresponding training stability metric (center), and the estimated spectrum of the local Jacobian in the complex plane (right). The spectral plot displays the eigenvalues computed from the block-Lanczos approximation of the empirical Jacobian, with the dashed unit circle indicating the boundary between contractive and divergent dynamics. The second row shows the evolution of the generator parameters (left) and discriminator parameters (center), represented by the Euclidean distance from their initialization, and the mode-collapse indicator (right). A nearly monotonic decrease of these distances together with a slow, smooth decay of the variance indicates stable and non-collapsing behavior. Overall, the combined diagnostics reveal that both LM-LRSGA variants maintain bounded losses, regular parameter trajectories, and eigenvalues concentrated within or near the unit circle, in contrast to the divergent and oscillatory dynamics observed with Adam. In particular, when using the LM–LRSGA–optimal configuration ($\eta=0.2$, $\tau=0.002$), Adam exhibits a spectral radius of $\rho=3.06$, a value well above the stability band $1-\varepsilon,1+\varepsilon]=[0.85,1.15]$ (with $\varepsilon=0.15$), and is therefore classified as unstable by the spectral criterion alone. For completeness, we also report the high–frequency (HF) power of the loss signals estimated via Welch’s method: both generator and discriminator HF powers are approximately of order $10^{3}$, confirming the presence of strong oscillatory energy.
In contrast, both LM–LRSGA and LM–LRSGA–EMA fall inside the tolerance band, with $\rho=1.1248$ and $\rho=1.1335$, respectively. For these borderline cases, we apply the Welch refinement: the orders of the HF power for generator and discriminator are $10^{-1}$ and $10^{0}$, respectively, for both LM–LRSGA and LM–LRSGA–EMA—values that are over three orders of magnitude smaller than for Adam and support their classification as marginally stable. Consistently, both limited–memory variants maintain nearly constant global stability indices ($\approx 0.99$) and parameter oscillations of order $\mathcal{O}(10^{-4})$, in line with their superior FID scores.
Under the Adam–optimal learning-rate configuration ($\eta=0.001$, $\tau=0.002$), LM–LRSGA and LM–LRSGA–EMA fall inside the tolerance band with $\rho\approx 1.06$ and $\rho\approx 1.06$, respectively, and are thus disambiguated via Welch’s PSD. Their high–frequency (HF) powers are of order $10^{-3}$ for both generator and discriminator, for both LM–LRSGA and LM–LRSGA–EMA, which are several orders of magnitude below those observed for Adam. Accordingly, both limited–memory variants are classified as marginally stable, with near–unit global stability indices and parameter oscillations essentially at $\mathcal{O}(10^{-4})$, indicating smooth, non–collapsing dynamics. In contrast, Adam attains $\rho\approx 1.48$ (above $1+\varepsilon=1.15$), hence is unstable by the spectral criterion alone; for completeness, its HF powers are of order $10^{-1}$ for the discriminator and $10^{1}$ for the generator, confirming residual oscillatory energy in the losses.
@@
Table 1: Spectral and stability metrics for Adam, LM-LRSGA, and LM-LRSGAEma under two learning-rate regimes. All quantities are computed from block-Lanczos spectral estimates and the auxiliary stability indices described in Section 7.1.
Setting Optimizer $\rho$ HF Power (G/D) Pred. Stab Stab. (G/D) Mode-Collapse
$\begin{array}[]{c}\eta=0.2\\ \tau=0.002\end{array}$ Adam 3.06 $1.1\times 10^{3}/1.0\times 10^{3}$ Unstable 0.63 / 0.78 High
LM-LRSGA 1.12 $7.1\times 10^{-1}/1.1\times 10^{0}$ Marginally Stable 0.99 / 0.99 Low
LM-LRSGAEma 1.13 $7.2\times 10^{-1}/1.1\times 10^{0}$ Marginally Stable 0.99 / 0.99 Low
$\begin{array}[]{c}\eta=0.001\\ \tau=0.002\end{array}$ Adam 1.48 $2.1\times 10^{1}/6.4\times 10^{-1}$ Unstable 0.99 / 0.98 Low
LM-LRSGA 1.06 $2.1\times 10^{-3}/3.5\times 10^{-3}$ Marginally Stable 1.00 / 1.00 Low
LM-LRSGAEma 1.06 $1.7\times 10^{-3}/3.2\times 10^{-3}$ Marginally Stable 1.00 / 1.00 Low
@@
Across both learning-rate regimes, the limited-memory variants substantially reduce the spectral radius compared with Adam and maintain marginally stable trajectories even for large step sizes. The reduction of $\rho$ and the absence of high-frequency energy in the loss spectra confirm that the LM-LRSGA correction effectively damps the antisymmetric component of the gradient field, yielding dynamics that are nearly contractive in the sense of the spectral criterion. These findings provide a consistent spectral explanation for the smoother convergence and improved sample quality obtained by LM-LRSGA in the FID experiments.
@@
(c) LM-LRSGAEma
Figure 4: Spectral and stability diagnostics for Adam, LM-LRSGA, and LM-LRSGAEma with learning rate $\eta=0.2$. Each panel shows (top) loss evolution in logarithmic scale, the training stability metric, and the estimated Jacobian spectrum in the complex plane; and (bottom) the generator and discriminator parameter evolution together with the mode-collapse indicator.
@@
(c) LM-LRSGAEma
Figure 5: Spectral and stability diagnostics for Adam, LM-LRSGA, and LM-LRSGAEma with learning rate $\eta=0.001$. Each panel shows (top) loss evolution in logarithmic scale, the training stability metric, and the estimated Jacobian spectrum in the complex plane; and (bottom) the generator and discriminator parameter evolution together with the mode-collapse indicator.
@@ 8 Conclusions
This manuscript extends our prior LRSGA work Vater et al. (2025) in four directions: (i) a new per-iteration spectral stability condition for LRSGA near Nash equilibria (Section 3); (ii) an $\varepsilon_{x},\varepsilon_{y}$-regularized mixed-block formulation that avoids forming the full Jacobian updates (Section 4); (iii) a limited-memory variant (LM-LRSGA) based on adapted recursions together with a local linear convergence analysis (Sections 4–5); (iv) an empirical study and spectral diagnostics in adversarial training (Sections 6–7). Our LM-LRSGA, building on symplectic gradient corrections and quasi-Newton approximations, retains the stabilizing effect of second-order information while reducing memory and compute via two-loop recursions algorithms. By storing only a short history of curvature pairs, it attains near–second-order stability with first-order–like efficiency.
Our analyses show that LM-LRSGA mitigates the rotational dynamics that typically hinder convergence in adversarial settings. On GAN training with MNIST and Fashion-MNIST, it yields smoother, synchronized loss trajectories, faster convergence, and FID improvement over Adam. These results indicate that limited-memory curvature corrections offer a practical balance of stability, efficiency, and scalability for competitive optimization.
Future work will focus on broadening the applicability of LM-LRSGA beyond two-player games. In particular, we plan to extend the method to competitive (or more generally non-cooperative) settings involving more than two neural architectures, and to assess its empirical and theoretical behavior in these multi-player training scenarios.
LM-LRSGA also supports sustainable, energy-aware machine learning. By avoiding explicit formation and storage of high-dimensional second-order tensors, it reduces memory transfers and expensive operationsthereby lowering runtime in line with “green AI” principles. Thus, LM-LRSGA provides a step toward methods that are not only theoretically and numerically robust, but also environmentally sustainable.